# An Efficient Row-Based Sparse Fine-Tuning

### v1

@@ 4.2.1 Class Aware Taylor Importance
${\bm{\eta}}_{i}^{t}:=|L(\mathcal{D}^{t},\acute{F}_{\bm{c}_{i}})-L(\mathcal{D}^{t},F)|\approx|{\mathbf{w}}^{\top}\nabla_{{\mathbf{w}}}L(\mathcal{D}^{t},F)|,$ (6)
where $F$ is the forward function of the entire model, $L(\mathcal{D}^{t},F)$ denotes the average loss of $F$ over inputs in class $t$, $\acute{F}_{\bm{c}_{i}}$ represents the forward without channel/neuron $\bm{c}_{i}$, and ${\mathbf{w}}$ is the parameter vector of channel $\bm{c}_{i}$. One natural choice of importance of neuron $i$ motivated by the above discussion is $\max_{t}{\bm{\eta}}_{i}^{t}$. We find that this score is “too sensitive” (importance of neurons may be over-estimated because of just one class), leading to lower overall accuracy. On the other hand, the average (over $t$) of ${\bf\eta}_{i}^{t}$ corresponds to the standard Taylor importance. We find that the intermediate quantity of Quantiles-Mean, defined as the average of the $0\%,10\%,20\%,\dots,100\%$ quantiles of the ${\bm{\eta}}_{i}^{t}$, works well in reducing the accuracy imbalance across labels, and also achieving a high overall accuracy. Formally,
@@ 4.2.2 Zero-Order Taylor Importance
As discussed, Taylor importance can incorporate information about the data distribution and the fine-tuning task when evaluating important neurons. However, for large models like Llama-3, it turns out that the computational overhead required for computing Taylor importances is prohibitively large! [^3] The Taylor importance here refers to computing the exact value without relying on approximations of the importance score or the gradient matrix used for deriving the importance score. In these cases, we apply the idea from the memory-efficient zeroth-order optimizer (MeZO, Malladi et al. 2023) to estimate the gradient in Taylor importance. The classical Zeroth-Order (ZO) setting is defined as below.
@@
Given a model $F$ with parameters $\mathbf{\theta}\in\mathbb{R}^{d}$ and a loss function $L$, SPSA estimates the gradient on a dataset $\mathcal{D}$ as
^$\hat{\nabla}L(\mathbf{\theta},\mathcal{D})=\frac{L(\mathbf{\theta}+\epsilon{\mathbf{z}})-L(\mathbf{\theta}-\epsilon{\mathbf{z}})}{2\epsilon}{\mathbf{z}},$
where ${\mathbf{z}}\in\mathbb{R}^{d}$ is drawn from ${\mathbf{z}}\sim\mathcal{N}(0,\mathbf{I}_{d})$, and $\epsilon$ is the scale of the perturbation. The $n$-SPSA estimate averages ^$\hat{\nabla}L(\mathbf{\theta},\mathcal{D})$ over $n$ randomly sampled ${\mathbf{z}}$. Note that as $\epsilon\rightarrow 0$, the estimate above converges to ${\mathbf{z}}({\mathbf{z}}^{\top}\nabla L(\mathbf{\theta},\mathcal{D}))$, which is equal to $\nabla L(\mathbf{\theta},\mathcal{D})$ in expectation.
@@
${\bm{\eta}}:=|\mathbf{\theta}^{\top}\hat{{\mathbf{g}}}|.$ (8)
where we denote $[\nabla L(\mathbf{\theta},\mathcal{D})]$ and its estimate as ${\mathbf{g}}$ and $\hat{{\mathbf{g}}}$ for convenience.
A naïve implementation of SPSA still requires twice the memory of inference because of the need to store ${\mathbf{z}}$. However, MeZO uses the trick of regenerating ${\mathbf{z}}$ dynamically using a random seed (of much smaller size than the model), thus eliminating the storage and ensuring memory usage that is equal to that of inference. We now assess the effectiveness of ZOTaylor for our LLM.
@@
$n$-SPSA is an unbiased and consistent estimator with the variance $\mathbf{\sigma}^{2}$ where
$\mathbf{\sigma}_{i}^{2}=\frac{{\mathbf{g}}_{i}^{2}+\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{n}.$
@@
Property 4.2 can be proved by simply noting that the covariance matrix of ${\mathbf{z}}$ is $\mathbf{I}_{d}$, details can be found in Appendix A. With this property, we know that $n$-SPSA can accurately estimate the gradient. However, the variance can be large when $d$ is large. Here, we first note that since we aim to find the most important neurons, we do not care about the gradient estimate itself. That is, given ${\mathbf{g}}_{i}>{\mathbf{g}}_{j}$, our goal is to have a higher probability $\Pr[\hat{{\mathbf{g}}}_{i}-\hat{{\mathbf{g}}}_{j}>0]$ which is exactly equal to:
@@
$\Pr_{Z\sim\mathcal{N}(0,1)}[Z>-\frac{{\mathbf{g}}_{i}-{\mathbf{g}}_{j}}{\sqrt{(\sigma_{i}^{2}+\sigma_{j}^{2})/2}}].$ (9)
@@
The lower bound of Equation 9 is $0.5$ and the probability will close to $0.5$ when the variance is large or when ${\mathbf{g}}_{i}$ and ${\mathbf{g}}_{j}$ are too close to one another. In our experiments, the values of $\hat{{\mathbf{g}}}$ from single SPSA lie in $[-100,100]$ with variances ranging from $10^{7}$ to $10^{8}$, thus the probability above turns out to be too close to $0.5$.
@@
To address this issue, we utilize a simple but highly effective strategy (J Reddi et al., 2015): we partition the training data into $k$ calibration sets. For each calibration set, we generate $n$ distinct perturbations ${\mathbf{z}}$ to perform $n$-SPSA, yielding $n\times k$ gradient estimates. Consequently, the variance of $n$-SPSA is effectively reduced to $\frac{{\mathbf{g}}_{i}^{2}+\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{nk}$. [^4] We note that this is not a formal guarantee; indeed, if $k$ is too large or the dataset is too small, there is additional variance across calibration sets that will become significant. In our experiments, we set $n=5$ and $k=256$, where the variances fall within the range of $[10^{4},10^{5}]$, leading to a normalized value $-\frac{{\mathbf{g}}_{i}-{\mathbf{g}}_{j}}{\sqrt{(\sigma_{i}^{2}+\sigma_{j}^{2})/2}}$ being approximately $-1$, which corresponds to the probability $\Pr[Z>-\frac{{\mathbf{g}}_{i}-{\mathbf{g}}_{j}}{\sqrt{(\sigma_{i}^{2}+\sigma_{j}^{2})/2}}]$ being close to $0.84$.
Importantly, our objective is to select most of the top-$x$important neurons (for appropriate $x$), rather than ensuring the strict selection of all top-ranked neurons. Thus, the probability $\Pr[\hat{{\mathbf{g}}}_{i}-\hat{{\mathbf{g}}}_{j}>0]$ for any specific $(i,j)$-pair does not need to be excessively high. A rigorous theoretical analysis of the optimal probability threshold is an interesting open direction. In this study, we primarily focus on demonstrating the feasibility of applying ZO-based approaches to importance metrics.
Model, ft setting mem #param ARC-c ARC-e BoolQ HS OBQA PIQA rte SIQA WG Avg
Llama2(7B)
LoRA, $r=64$ 23.46GB 159.9M(2.37%) 44.97 77.02 77.43 57.75 32.0 78.45 62.09 47.75 68.75 60.69
VeRA, $r=64$ 22.97GB 1.374M(0.02%) 43.26 76.43 77.40 57.26 31.6 78.02 62.09 45.85 68.75 60.07
DoRA, $r=64$ 44.85GB 161.3M(2.39%) 44.71 77.02 77.55 57.79 32.4 78.29 61.73 47.90 68.98 60.71
RoSA, $r=32,d=1.2\%$ 44.69GB 157.7M(2.34%) 43.86 77.48 77.86 57.42 32.2 77.97 63.90 47.29 69.06 60.78
SPruFT, $r=128$ 17.62GB 145.8M(2.16%) 43.60 77.26 77.77 57.47 32.6 78.07 64.98 46.67 69.30 60.86
Llama3(8B)
LoRA, $r=64$ 30.37GB 167.8M(2.09%) 53.07 81.40 82.32 60.67 34.2 79.98 69.68 48.52 73.56 64.82
VeRA, $r=64$ 29.49GB 1.391M(0.02%) 50.26 80.30 81.41 60.16 34.4 79.60 69.31 46.93 72.77 63.90
DoRA, $r=64$ 51.45GB 169.1M(2.11%) 53.33 81.57 82.45 60.71 34.2 80.09 69.31 48.67 73.64 64.88
RoSA, $r=32,d=1.2\%$ 48.40GB 167.6M(2.09%) 51.28 81.27 81.80 60.18 34.4 79.87 69.31 47.95 73.16 64.36
SPruFT, $r=128$ 24.49GB 159.4M(1.98%) 52.47 81.10 81.28 60.29 34.6 79.76 70.04 47.75 73.24 64.50
Table 3: Fine-tuning Llama on Alpaca dataset for 5 epochs and evaluating on 9 tasks from EleutherAI LM Harness. “mem” represents the memory usage, with further details provided in Appendix C.3. #param is the number of trainable parameters, where the difference of #param between the two approaches depends on the architecture of Llama, as some layers have $d_{in}\neq d_{out}$. HS, OBQA, and WG represent HellaSwag, OpenBookQA, and WinoGrande datasets. The ablation study for different $r$ can be found in Appendix C.6. All reported results are accuracies on the corresponding tasks. Bold indicates the best result on the same task.
@@ 5 Experimental Setup
5.1 Datasets
We use multiple datasets for different tasks. For image classification, we fine-tune models on the training split and evaluate it on the validation split of Tiny-ImageNet (Tavanaei, 2020), CIFAR100 (Krizhevsky et al., 2009), and Caltech101 (Li et al., 2022). For text generation, we fine-tune LLMs on 256 samples from Stanford-Alpaca (Taori et al., 2023) and assess zero-shot performance on nine EleutherAI LM Harness tasks (Gao et al., 2021). See Appendix D for details.
5.2 Models and Baselines
@@
We fine-tune full-precision Llama-2-7B and Llama-3-8B (float32) using our SPruFT, LoRA (Hu et al., 2022), VeRA (Kopiczko et al., 2024), DoRA (Liu et al., 2024), and RoSA (Nikdan et al., 2024). RoSA is chosen as the representative SFT method and is the only SFT due to the high memory demands of other SFT approaches, while full fine-tuning is excluded for the same reason. We freeze Llama’s classification layers and fine-tune only the linear layers in attention and MLP blocks.
Next, we evaluate importance metrics by fine-tuning Llamas and image models, including DeiT (Touvron et al., 2021), ViT (Dosovitskiy, 2020), ResNet101 (He et al., 2016), and ResNeXt101 (Xie et al., 2017) on CIFAR100, Caltech101, and Tiny-ImageNet. For image tasks, we set the fine-tuning ratio at 5%, meaning the trainable parameters are a total of 5% of the backbone plus classification layers.
5.3 Training Details
@@
Our fine-tuning framework is built on torch-pruning [^5] Torch-pruning is not required, all their implementations are based on PyTorch. (Fang et al., 2023), PyTorch (Paszke et al., 2019), PyTorch-Image-Models (Wightman, 2019), and HuggingFace Transformers (Wolf et al., 2020). Most experiments run on a single A100-80GB GPU, while DoRA and RoSA use an H100-96GB GPU. We use the Adam optimizer (Kingma & Ba, 2015) and fine-tune all models for a fixed number of epochs without validation-based model selection.
Image models: The learning rate is set to $10^{-4}$ with cosine annealing decay (Loshchilov & Hutter, 2017), where the minimum learning rate is $10^{-9}$. All image models used in this study are pre-trained on ImageNet.
Llama: For LoRA and DoRA, we set $\alpha=16$, a dropout rate of $0.1$, and a learning rate of $10^{-4}$ with linear decay ( $0.01$ decay rate). For SPruFT, we control trainable parameters using rank instead of fine-tuning ratio for direct comparison. The learning rate is $2\cdot 10^{-5}$ with the same decay settings. Linear decay is applied after a warmup over the first $3$% of training steps. The maximum sequence length is $2048$, with truncation for longer inputs and padding for shorter ones.

### v2

@@ 4.3.1 Class Aware Taylor Importance
${\bm{\eta}}_{i}^{t}:=|L(\mathcal{D}^{t},\acute{F}_{i})-L(\mathcal{D}^{t},F)|\approx|({\bm{\theta}}^{(i)})^{\top}\nabla_{{\bm{\theta}}^{(i)}}L(\mathcal{D}^{t},F)|,$ (7)
where $F$ is the forward function of the entire model, $L(\mathcal{D}^{t},F)$ denotes the average loss of $F$ over inputs in class $t$, $\acute{F}_{i}$ represents the forward without channel/neuron $i$, and ${\bm{\theta}}^{(i)}$ is the parameter vector of channel $i$. One natural choice of importance of neuron $i$ motivated by the above discussion is $\max_{t}{\bm{\eta}}_{i}^{t}$. We find that this score is “too sensitive” (importance of neurons may be over-estimated because of just one class), leading to lower overall accuracy. On the other hand, the average (over $t$) of ${\bf\eta}_{i}^{t}$ corresponds to the standard Taylor importance. We find that the intermediate quantity of Quantiles-Mean, defined as the average of the $0\%,10\%,20\%,\dots,100\%$ quantiles of the ${\bm{\eta}}_{i}^{t}$, works well in reducing the accuracy imbalance across labels, and also achieving a high overall accuracy. Formally,
@@ 4.3.2 Zero-Order Taylor Importance
As discussed, Taylor importance can incorporate information about the data distribution and the fine-tuning task when evaluating important neurons. However, for large models like Llama-3, it turns out that the computational overhead required for computing Taylor importances is prohibitively large. [^3] The Taylor importance here refers to computing the exact value without relying on approximations of the importance score or the gradient matrix used for deriving the importance score. In these cases, we apply the idea from the memory-efficient zeroth-order optimizer (MeZO, Malladi et al. 2023) to estimate the gradient in Taylor importance. The classical Zeroth-Order (ZO) setting is defined as below.
@@
Given a model with parameters ${\bm{\theta}}\in\mathbb{R}^{d}$ and a loss function $L$, SPSA estimates the gradient on a dataset $\mathcal{D}$ as
^$\hat{\nabla}L({\bm{\theta}},\mathcal{D})=\frac{L({\bm{\theta}}+\epsilon{\mathbf{z}})-L({\bm{\theta}}-\epsilon{\mathbf{z}})}{2\epsilon}{\mathbf{z}},$
where ${\mathbf{z}}\in\mathbb{R}^{d}$ is drawn from ${\mathbf{z}}\sim\mathcal{N}(0,\mathbf{I}_{d})$, and $\epsilon$ is the scale of the perturbation. The $n$-SPSA estimate averages ^$\hat{\nabla}L({\bm{\theta}},\mathcal{D})$ over $n$ randomly sampled ${\mathbf{z}}$. Note that as $\epsilon\rightarrow 0$, the estimate above converges to ${\mathbf{z}}({\mathbf{z}}^{\top}\nabla L({\bm{\theta}},\mathcal{D}))$, which is equal to $\nabla L({\bm{\theta}},\mathcal{D})$ in expectation.
@@
${\bm{\eta}}_{i}:=|({\bm{\theta}}^{(i)})^{\top}{\mathbf{g}}^{(i)}|,\hat{\bm{\eta}}_{i}:=|({\bm{\theta}}^{(i)})^{\top}\hat{{\mathbf{g}}}^{(i)}|,$ (9)
where we denote $[\nabla L({\bm{\theta}},\mathcal{D})]$ and its estimate as ${\mathbf{g}}$ and $\hat{{\mathbf{g}}}$ for convenience, and ${\bm{\eta}}_{i}$, $({\bm{\theta}}^{(i)})^{\top}$, and ${\mathbf{g}}^{(i)}$ are the importance score, the parameter vector, and the gradient vector for neuron $i$.
We now assess the efficiency and effectiveness of ZOTaylor for our LLM. For efficiency, a naïve implementation of SPSA still requires twice the memory of inference because of the need to store ${\mathbf{z}}$. However, MeZO uses the trick of regenerating ${\mathbf{z}}$ dynamically using a random seed (of much smaller size than the model), thus eliminating the storage and ensuring memory usage that is equal to that of inference. We then justify the effectiveness of ZOTaylor in the following.
@@
$n$-SPSA is an unbiased and consistent estimator with the variance vector $({\sigma}_{1}^{2},\cdots,{\sigma}_{d}^{2})$ where
$\mathbf{\sigma}_{j}^{2}=\frac{{\mathbf{g}}_{j}^{2}+\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{n}.$
@@
Property 4.2 can be proved by simply noting that the covariance matrix of ${\mathbf{z}}$ is $\mathbf{I}_{d}$, details can be found in Appendix B. With this property, we know that $n$-SPSA can accurately estimate the gradient. However, the variance can be large when $d$ is large. Here, we first note that since we aim to find the most important neurons, we do not care about the gradient estimate itself. That is, given ${\bm{\eta}}_{i_{1}}>{\bm{\eta}}_{i_{2}}$, our goal is to have a higher probability $p\left(\hat{\bm{\eta}}_{i_{1}}-\hat{\bm{\eta}}_{i_{2}}>0\right)$ which is exactly equal to:
@@
$\displaystyle p_{Z\sim\mathcal{N}(0,1)}\left(Z>-\frac{|{\mathbf{g}}^{(i_{1})}{\bm{\theta}}^{(i_{1})}|-|{\mathbf{g}}^{(i_{2})}{\bm{\theta}}^{(i_{2})}|}{\sqrt{\left(\frac{\sum_{j\in\{j^{(i_{1})}\}}\mathbf{\sigma}_{j}^{2}{\bm{\theta}}_{j}^{2}}{2}+\frac{\sum_{j\in\{j^{(i_{2})}\}}\mathbf{\sigma}_{j}^{2}{\bm{\theta}}_{j}^{2}}{2}\right)/2}}\right),$ (10)
@@
where $\{j^{(i_{1})}\}$ and $\{j^{(i_{2})}\}$ are the indices of neuron $i_{1}$ and $i_{2}$, and $d_{i_{1}}=|\{j^{(i_{1})}\}|,d_{i_{2}}=|\{j^{(i_{2})}\}|$. The trivial lower bound of Equation 10 is $0.5$ and the probability will close to $0.5$ when the variance is large or when ${\bm{\eta}}_{i_{1}}$ and ${\bm{\eta}}_{i_{2}}$ are too close to one another. In our experiments, the values of $\hat{{\mathbf{g}}}_{j}$ from a single SPSA are almost uniformly in $[-100,100]$ with variances ranging from $10^{7}$ to $10^{8}$ and the values of ${\bm{\theta}}_{j}$ are mostly in $[-1,1]$, thus the probability above turns out to be too close to $0.5$.
@@
To address this issue, we utilize a simple but highly effective strategy (J Reddi et al., 2015): we partition the training data into $k$ calibration sets. For each calibration set, we generate $n$ distinct perturbations ${\mathbf{z}}$ to perform $n$-SPSA, producing $n\times k$ gradient estimates. Consequently, the variance of $n$-SPSA is effectively reduced to $\frac{{\mathbf{g}}_{j}^{2}+\sum_{l=1}^{d}{\mathbf{g}}_{l}^{2}}{nk}$ [^4] We note that this is not a formal guarantee; indeed, if $k$ is too large or the dataset is too small, there is additional variance across calibration sets that will become significant..
Theorem 4.3.
Suppose a model parameter ${\bm{\theta}}$ satisfies $|{\bm{\theta}}|_{\infty}\leq u_{\bm{\theta}}$ and a gradient ${\mathbf{g}}$ satisfies $|{\mathbf{g}}|_{2}\leq u_{\bm{{\mathbf{g}}}}$ for some parameter $u_{\bm{{\mathbf{g}}}}$. Also assume that ${\bm{\eta}}_{i}$’s are drawn IID from a distribution $\mathcal{P}_{\bm{\eta}}$ that is $\alpha$-smooth and supported in the interval $[0,u_{\bm{\eta}}]$. Let ${\bm{\eta}}_{i_{1}}$ and ${\bm{\eta}}_{i_{2}}$ be the importance values of the top-$q_{1}$and top-$q_{2}$important neurons and define $\Phi$ to be the CDF of Gaussian distribution, then we have
$\displaystyle p\left(\hat{\bm{\eta}}_{i_{1}}-\hat{\bm{\eta}}_{i_{2}}>0\right)\geq\displaystyle p_{Z\sim\mathcal{N}(0,1)}\left(Z>-\frac{2\sqrt{nk}({\bm{\eta}}_{i_{1}}-{\bm{\eta}}_{i_{2}})}{\sqrt{(d_{i_{1}}+d_{i_{2}})u_{\bm{\theta}}^{2}u_{\mathbf{g}}^{2}}}\right)\geq 1-\Phi\left(-\frac{2\sqrt{nk}(q_{2}-q_{1})/100}{\alpha\sqrt{(d_{i_{1}}+d_{i_{2}})u_{\bm{\theta}}^{2}u_{\mathbf{g}}^{2}}}\right).$ (11)
Theorem 4.3 follows directly from the variance expression $\mathbf{\sigma}_{j}$ provided in Property 4.2; detailed steps are included in Appendix B. Using this result, we now show how to estimate the required value of $\sqrt{nk}$ to reliably identify most of the desired neurons.
We argue that for any $\xi$, if we select the top-$\xi$important neurons according to the $\hat{\bm{\eta}}$ values (for some $\xi\in[0,100]$), most of top-$\xi$neurons according to the true ${\bm{\eta}}$ values will be correctly selected with high probability, when $nk$ is large enough. To formalize this, let $i_{\xi}=\frac{\xi}{100}d_{\bm{\eta}}$ where $d_{\bm{\eta}}$ is the number of neurons and let ${\bm{\eta}}_{1}\geq{\bm{\eta}}_{2}\geq\ldots\geq{\bm{\eta}}_{d_{\bm{\eta}}}$. Let $X_{i}$ be the random variable that is $1$ if $\hat{\bm{\eta}}_{i}$ is among the top $i_{\xi}$ of $\{\hat{\bm{\eta}}_{j}\}_{j\in[d_{\bm{\eta}}]}$ and 00 otherwise. Define $p_{i}$ to be $p\left(X_{i}=1\right)$.
We can now bound $p(X_{i}=0)$ as follows. If $X_{i}=0$ for $i=i_{\xi}$, then at least one of the values $\hat{\bm{\eta}}_{i_{\xi}+1},\hat{\bm{\eta}}_{i_{\xi}+2},\ldots,\hat{\bm{\eta}}_{d_{\bm{\eta}}}$, say $\hat{\bm{\eta}}_{j}$, must be $>\hat{\bm{\eta}}_{i}$. Thus, for all $i\leq i_{\xi}$
$p(X_{i}=0)\leq 1-\prod_{j=i_{\xi}+1}^{d_{\bm{\eta}}}\displaystyle p(\hat{\bm{\eta}}_{i}>\hat{\bm{\eta}}_{j})\leq 1-\displaystyle p(\hat{\bm{\eta}}_{i}>\hat{\bm{\eta}}_{i_{\xi}})^{d_{\bm{\eta}}-i_{\xi}}\implies\displaystyle p_{i}\geq\displaystyle p(\hat{\bm{\eta}}_{i}>\hat{\bm{\eta}}_{i_{\xi}})^{d_{\bm{\eta}}-i_{\xi}}.$ (12)
The next result bounds the expected number of the “correct” neurons (or indices) chosen, in terms of $n,k$.
Theorem 4.4.
Let $\epsilon_{\xi}\in(0,1]$ be any desired error ratio, and let $X_{i}$ be as defined above. Then there exist a positive constant $c_{d_{\bm{\eta}}}$ such that the inequality $\mathbb{E}[\sum_{i=1}^{i_{\xi}}X_{i}]\geq(1-\epsilon_{\xi})\cdot i_{\xi}$ holds, as long as
$\sqrt{nk}\geq c_{d_{\bm{\eta}}}\cdot\alpha\sqrt{(d_{i_{1}}+d_{i_{2}})u_{\bm{\theta}}^{2}u_{\mathbf{g}}^{2}}.$ (13)
Theorem 4.4 can be proved by using Equation 12 and Theorem 4.3; detailed steps are given in Appendix B.
Unfortunately, the above threshold of $\sqrt{nk}$ is too large in our experiments. To balance the variance of estimates and computation resources, we set $nk=2048$ in our experiments. The error $\epsilon_{\xi=10}$ is likely to be higher than $20$% under this setting, but our experimental results show that this setting is good enough to boost the fine-tuning performance.
@@ 5 Experimental Setup
Main experiments:
For language tasks, we fine-tune LLaMA-2-7B and LLaMA-3-8B in both full-precision (float32) and NF4 (Dettmers et al., 2024) on the training splits of 8 commonsense reasoning datasets (see Appendix E), and evaluate them on their respective test splits. We then assess the models’ mathematical reasoning by fine-tuning them on the GSM8K (Cobbe et al., 2021) training split and evaluating performance using the EleutherAI LM Harness (Gao et al., 2021). Finally, we evaluate instruction-following ability by fine-tuning on Alpaca-GPT (Taori et al., 2023) and scoring responses with MT-Bench (Zheng et al., 2023), using Gemini 2.5 flash (Team et al., 2023) as the judge [^5] We use Gemini 2.5 instead of GPT-4 due to lower cost and competitive performance..
@@
We incorporate the PEFT approaches including our SPruFT, QSPruFT, LoRA (Hu et al., 2022), QLoRA (Dettmers et al., 2024), VeRA (Kopiczko et al., 2024), DoRA (Liu et al., 2024), QDoRA, RoSA (Nikdan et al., 2024), LoftQ (Li et al., 2024), and QPiSSA (Meng et al., 2024). RoSA is chosen as the representative SpFT method and is the only SpFT due to the high memory demands of other SpFT approaches, while full fine-tuning is excluded for the same reason. We freeze Llama’s classification layers and fine-tune only the linear layers in attention and FFN blocks.
For training configurations setting, we use a learning rate of $2\cdot 10^{-5}$ with linear decay (rate 0.01) for our method, and $10^{-4}$ for other PEFT methods, with $\alpha=16$ and dropout rate $0.1$. All methods apply linear decay after a 3% warmup. For commonsense reasoning, we train on 2048 samples (256 per dataset) and evaluate on 500 test examples per dataset. For GSM8K, we fine-tune on 2048 random training samples and evaluate on the full test set. For Alpaca-GPT, we train the models on 1024 samples and evaluate the models on the whole MT-bench. Instruction-style prompts are used for commonsense datasets, [^6] LLaMA-3 performs well with question-answering prompts, and fine-tuning yields limited gains, suggesting possible pre-training on these datasets with question-answering prompts. while GSM8K uses question-answering prompts. We fine-tune all models for 3 epochs.
@@
Our framework is built on torch-pruning (Fang et al., 2023), PyTorch (Paszke et al., 2019), and HuggingFace Transformers (Wolf et al., 2020). Most experiments are conducted on a single A100-80GB GPU, except DoRA and RoSA (at 2048 max tokens) which run on an H100-96GB. We use the Adam optimizer (Kingma & Ba, 2015) and train with a fixed epoch budget without early stopping.
Experiments for importance metrics: In the comparison of different importance metrics, we also use our approach to fine-tune DeiT (Touvron et al., 2021), ViT (Dosovitskiy, 2020), ResNet101 (He et al., 2016), and ResNeXt101 (Xie et al., 2017) on Tiny-ImageNet (Tavanaei, 2020), CIFAR100 (Krizhevsky et al., 2009), and Caltech101 (Li et al., 2022). For these tasks, we set the fine-tuning ratio at 5%, meaning the trainable parameters are a total of 5% of the backbone plus classification layers. Following this, we discuss the computational resource requirements for fine-tuning.
For image models, the learning rate is set to $10^{-4}$ with cosine annealing decay (Loshchilov & Hutter, 2017), where the minimum learning rate is $10^{-9}$. All image models used in this study have been pre-trained on ImageNet. Note that memory efficiency is not emphasized for small-scale models, as dataset-related memory—particularly with large batch sizes—dominates consumption in these cases. The main advantage of our method in these cases is the reduced FLOPs due to fewer trainable parameters.
Table 2: Main results of fine-tuning full precision Llama2 and Llama3. “mem” represents the memory cost in training excluding the model itself, with further details provided in Appendix D.3. #param is the number of trainable parameters. HS, OBQA, and WG represent HellaSwag, OpenBookQA, and WinoGrande datasets. All reported results for commonsense reasoning tasks are accuracies, while the results for GSM8k are the exact match score. The ablation study for different $r$ can be found in Appendix D.6. All reported results are accuracies on the corresponding tasks. Bold indicates the best result on the same task.
Model, ft setting mem #param BoolQ PIQA SIQA HS WG ARC-c ARC-e OBQA Avg GSM8k
Llama2(7B), pretrained - - 58.00 40.00 29.00 15.40 4.80 14.00 16.20 25.40 25.35 0.00
LoRA, $r=64$ 23.46GB 159.9M(2.37%) 77.00 76.20 67.80 84.20 62.60 70.00 82.00 74.00 74.23 18.42
VeRA, $r=64$ 22.97GB 1.374M(0.02%) 47.80 51.80 41.80 37.60 50.40 36.80 43.20 32.80 41.53 0.00
DoRA, $r=64$ 44.85GB 161.3M(2.39%) 75.20 75.40 64.60 78.60 63.00 65.20 82.20 70.60 71.85 21.46
RoSA, $r=32,d=1.2\%$ 44.69GB 157.7M(2.34%) 79.80 73.40 70.20 76.00 57.00 68.80 80.80 71.60 72.20 21.99
SPruFT, $r=128$ 17.62GB 145.8M(2.16%) 80.00 75.20 67.60 85.00 63.40 70.80 82.40 71.80 74.53 22.90
Llama3(8B), pretrained - - 58.80 41.60 38.00 10.20 11.20 55.20 63.00 27.40 38.18 0.00
LoRA, $r=64$ 30.37GB 167.8M(2.09%) 84.20 77.00 63.20 84.20 67.20 76.40 88.80 71.00 76.50 41.77
VeRA, $r=64$ 29.49GB 1.391M(0.02%) 61.00 62.40 55.60 41.80 49.60 59.60 77.60 60.00 58.45 0.00
DoRA, $r=64$ 51.45GB 169.1M(2.11%) 83.20 82.80 69.00 89.40 70.80 77.20 89.00 80.40 80.23 46.02
RoSA, $r=32,d=1.2\%$ 48.40GB 167.6M(2.09%) 79.00 81.00 69.20 84.80 68.60 79.00 90.40 78.40 78.80 45.72
SPruFT, $r=128$ 24.49GB 159.4M(1.98%) 87.60 77.40 71.40 85.40 70.20 79.80 90.80 81.80 80.55 46.10